### v2

@@ Abstract
This paper proposes a nonparametric test of pairwise independence of one random variable from a large pool of other random variables. The test statistic is the maximum of several Chatterjee’s rank correlations and critical values are computed via a block multiplier bootstrap. The test is shown to asymptotically control size uniformly over a large class of data-generating processes, even when the number of variables is much larger than sample size. The test is consistent against any fixed alternative. It can be combined with a stepwise procedure for selecting those variables from the pool that violate independence, while controlling the family-wise error rate. All formal results leave the dependence among variables in the pool completely unrestricted. In simulations, we find that our test is very powerful, outperforming existing tests in most scenarios considered, particularly in high dimensions and/or when the variables in the pool are dependent.
Keywords: independence test, high-dimensional data, Chatterjee’s rank correlation, block multiplier bootstrap, family-wise error rate.
@@ 1 Introduction
While it has been shown that the nonparametric bootstrap does not consistently estimate the distribution of a single Chatterjee’s rank correlation (Lin and Han, 2024), the proposed block multiplier bootstrap achieves this by accounting for the dependence of the summands in the asymptotically linear representation. The existence of an asymptotic linear representation for Chatterjee’s rank correlation together with suitably strong control of the remainder terms allows us to employ recent results for the approximation of maxima of high-dimensional vectors of sums (Chernozhukov, Chetverikov, and Kato, 2013, 2015, 2019) to establish the validity of the block multiplier bootstrap also for the maximum of many Chatterjee’s rank correlation coefficients.
The block multiplier bootstrap requires a tuning parameter choice, namely the block size that should tend to infinity with sample size. We are able to provide a simple choice that enjoys a certain optimality property: it minimizes the distance of the bootstrap variance and the asymptotic variance for each individual test statistic. In simulations, however, we find that the size and power of our test are almost insensitive to the particular value of the block size.
@@
Our paper is most closely related to the literature on testing independence of two random vectors, $H_{0}^{joint}\colon Y\perp X$, where $Y\in\mathbb{R}^{p}$ and $X\in\mathbb{R}^{q}$. This is a different hypothesis from the one we consider; for $q=1$, it implies, but is not implied by, $H_{0}$. Therefore, tests that control the rejection probability under $H_{0}^{joint}$ are not guaranteed to control it under our hypothesis $H_{0}$. Sinha and Wieand (1977), Taskinen, Oja, and Randles (2005), Bakirov, Rizzo, and Székely (2006), Székely, Rizzo, and Bakirov (2007), Heller, Gorfine, and Heller (2012), Heller, Heller, and Gorfine (2012), Shi, Hallin, Drton, and Han (2022) propose nonparametric tests of $H_{0}^{joint}$, where $p$ and $q$ are of arbitrary, but fixed (with sample size) dimensions. Székely and Rizzo (2013) show that the test statistic of Székely, Rizzo, and Bakirov (2007) is biased in high dimensions and an independence test based on it therefore does not control size in high dimensions. They also propose a bias-corrected test statistic and derive its asymptotic distribution under the null of independence when the dimensions of both vectors grow with the sample size. The asymptotic regime under which their test is valid requires the dimensions of both vectors to grow, so it is not clear (at least to us) whether it is also valid when one of the two dimensions remains constant as the sample size grows. In addition, their derivation of the test statistic’s limiting distribution requires the elements of the two vectors to be exchangeable, a condition we do not require for $Y_{1},\ldots,Y_{p}$. Ramdas, Jakkam Reddi, Poczos, Singh, and Wasserman (2015) show that both independence tests, Székely, Rizzo, and Bakirov (2007) and Székely and Rizzo (2013) have low power against “fair alternatives” in high dimensions. More recently, Zhou, Xu, Zhu, and Li (2024) and Wang, Liu, and Feng (2024) propose other rank-based tests, e.g. based on Hoeffding’s D, Blum-Kiefer-Rosenblatt’s R and Bergsma-Dassios-Yanagimoto’s $\tau$ among others, of $H_{0}^{joint}$ in high dimensions. The validity of these tests relies on at least one of the dimensions $p$ and/or $q$ diverging so that a central limit theorem across the elements of, say, $Y$ can be invoked. This approach necessarily restricts the dependence of $Y_{1},\ldots,Y_{p}$, while our validity results leave the dependence completely unrestricted. In simulations, we compare our test to the distance covariance tests by Székely, Rizzo, and Bakirov (2007), Székely and Rizzo (2013), Zhu, Zhang, Yao, and Shao (2020), and various tests proposed in Zhou, Xu, Zhu, and Li (2024). Even though we are not aware of theoretical guarantees that these tests of $H_{0}^{joint}$ also control the rejection probability under our null, $H_{0}$, in the simulations, we find that they do. However, in most scenarios considered, our test is substantially more powerful than all the other tests, particularly in high dimensions and/or when the variables $Y_{1},\ldots,Y_{p}$ are dependent. An additional contribution of our paper to this literature is our proposal of a stepwise procedure for selecting hypotheses that are not rejected so as to control the family-wise error rate.
@@
There is a large literature on nonparametric tests of mutual independence among the elements of a random vector. Some examples are Leung and Drton (2018), Yao, Zhang, and Shao (2018), Drton, Han, and Shi (2020), Wang, Liu, Feng, and Ma (2024), Bastian, Dette, and Heiny (2024); see also references therein. Xia, Cao, Du, and Dai (2024) propose such a test based on Chatterjee’s rank correlation. While the hypothesis considered in our paper also involves many nonparametric independence tests, it substantially differs from the null of mutual independence. This is because, in our testing problem, $X$ occurs in every independence hypothesis and the dependence of $Y_{1},\ldots,Y_{p}$ is left unrestricted.
@@
Finally, since our proposed test is based on the rank correlation for two random variables proposed by Chatterjee (2021), our paper is also related to a recent and fast-growing literature that examines the rank correlation coefficient’s properties. Chatterjee (2021) shows asymptotic normality of the correlation coefficient under independence of the two random variables. Lin and Han (2022a) and Kroll (2024) show that it is also asymptotically normal under dependence. Shi, Drton, and Han (2021) and Lin and Han (2022b) examine and propose improvements of the power of tests of independence based on the rank correlation coefficient. Based on the earlier work by Chatterjee (2021), Azadkia and Chatterjee (2021) introduce a graph-based correlation coefficient that can be seen as a multivariate extension of Chatterjee’s correlation coefficient. Han and Huang (2024) derive its asymptotic distribution under the null that a vector $Y$ (with fixed dimension) is independent of a random variable $X$, and they find, in simulations, that the test may be more powerful than competitors in higher dimensions. For a recent review of this literature, see Chatterjee (2024).
@@ 2 The Test
$\hat{\xi}_{j}\coloneqq 1-\frac{3n}{n^{2}-1}\sum_{i=1}^{n-1}\left\lvert\hat{F}_{Y_{j}}(Y_{j,(i+1)})-\hat{F}_{Y_{j}}(Y_{j,(i)})\right\rvert\leavevmode\nobreak\ .$ (1)
@@
[…]
a measure of dependence introduced by Dette, Siburg, and Stoimenov (2013) in the case in which $Y_{j}$ has a continuous distribution with density $f_{Y_{j}}$. This measure has several attractive features (Chatterjee, 2021). Two features that are particularly important for the test proposed in this paper are that (i) $\xi_{j}$ is equal to zero if, and only if, $X$ and $Y_{j}$ are independent, and (ii) the estimator $\hat{\xi}_{j}$ admits an asymptotic linear representation (shown in (3) below) with a remainder that we can show to be sufficiently small. Our arguments for validity of the proposed test in high dimensions crucially depend on property (ii).
@@
$\sqrt{n}\,\hat{\xi}_{j}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}+r_{j,n}\leavevmode\nobreak\ ,$ (3)
@@
with $U_{j,i}\coloneqq F_{Y_{j}}(Y_{j,(i)})$. A naive application of the multiplier bootstrap idea would be to repeatedly draw bootstrap multipliers $\varepsilon_{1},\ldots,\varepsilon_{n}$ as independent standard normal random variables that are independent of the data $\mathbb{D}$ and then compute a critical value from the distribution of $\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}\varepsilon_{i}W_{j,i}$ given the data. However, there are two problems with this approach. First, $\{W_{j,i}\}_{i=1}^{n}$ is not an i.i.d. sequence but a 1-dependent process. Second, $\{W_{j,i}\}_{i=1}^{n}$ are not directly observed because $F_{Y_{j}}$ is unknown.
To address the first challenge, we decompose the sum $\sum_{i=1}^{n-1}W_{j,i}$ into the sums over “big” and “small” blocks formed of $\{W_{j,i}\}_{i=1}^{n}$ with the property that the big blocks are independent of each other. Formally, let $q\geq 1$ denote the size of big blocks. It is a tuning parameter to be chosen by the researcher; in Section 2.3 we develop an optimal choice of $q$. Since $\{W_{j,i}\}_{i=1}^{n}$ is a 1-dependent sequence, we let the small blocks be of length one. Further, let $m\coloneqq\lfloor(n-1)/(q+1)\rfloor$, where $\lfloor\nu\rfloor$ is the integer part of $\nu$, denote the number of big blocks. Lastly, $r\coloneqq n-1-m(q+1)$, $10\leq r<q+$ 1, is the number of remaining summands that are not allocated to any of the small or big blocks. With this notation, we can thus write
@@ 2.1 Size Control
Suppose $p\geq 2$. There exist constants $C_{1}>0$ and $40<\gamma<1/$ 4 such that $(1/q)\log^{2}p\leq C_{1}n^{-\gamma}$ and $\max\big{\{}q\log^{5/2}p,\,\sqrt{q}\log^{7/2}(pn)\big{\}}\leq C_{1}n^{1/2-\gamma}$.
@@
$T_{0}\coloneqq\max_{1\leq j\leq p}\,\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}\leavevmode\nobreak\ ,$
@@
Step (iii) of the derivation combines several results on high-dimensional Gaussian approximations from Chernozhukov, Chetverikov, and Kato (2013, 2015, 2019). The main challenge of the proof is contained in steps (i) and (ii), where we need to establishing that all remainder terms $r_{1,n},\ldots,r_{p,n}$ from the representation (3) vanish at a suitably fast rate. Angus (1995) shows that, for each $j$, $r_{j,n}=o_{P}(n^{-1/2})$, but in our high-dimensional setting, we need stronger control of these remainder terms, and these results are developed in Lemma 2.
@@
Remark 3 (studentization).
The test proposed in this section does not studentize the test statistics. In simulations in Section 4, we find that studentization may improve size and power of the test. We consider studentizing the individual test statistics by their standard deviation under the null, i.e.,
@@
For the bootstrap statistics there are at least two different possibilities for studentization. First, one could studentize it by the square root of $\operatorname*{\mathbb{E}}[A_{j,k}^{2}]/q=0.4+0.1/q$, i.e.
[…]
This standardization ensures that the diagonal elements of the bootstrap covariance matrix of individual tests are all approximately centered at one for any $q$, and hence approximately match the variance of $\sqrt{n}\hat{\xi}_{j}/\sqrt{v_{n}}$. The second possibility is to employ a bootstrap test statistic in which the big blocks sums are demeaned and standardized by their sample standard deviation, i.e.
[…]
This standardization ensures that the diagonal elements of the bootstrap covariance matrix of individual test statistics are all equal to one for any $q$.
@@ 2.2 Consistency
In light of recent discussions related to the local asymptotic power of Chatterjee’s test of independence (e.g., Shi, Drton, and Han (2021) and Lin and Han (2022b)), it would be interesting to study the power of our proposed test to local alternatives. We view this analysis as beyond the scope of the paper and relegate it to future research.
@@ 2.3 Choice of $q$
In our bootstrap procedure, $T^{B}_{0,j}$ mimics the behavior of the individual test statistic $\sqrt{n}\hat{\xi}_{j}$, which is asymptotically normally distributed and, under the null, has variance $v_{n}$ given in Remark 3. Since $q$ affects the conditional distribution of $T^{B}_{0,j}$ only through the bootstrap variance $V^{B}_{j}$ and it does not affect the test statistic itself, we aim to choose $q$ such that $V^{B}_{j}$ is close to $v_{n}$.
@@
Minimizing $MSE_{j,n}(q)$ over $q\in\mathbb{N}^{+}$ yields our proposed optimal choice of $q$:
@@
The minimizer $q^{*}(n)$ can be computed for each $n$ by evaluating the function $MSE_{j,n}(q)$ over a grid of values $q\in\{1,2,\ldots,\lfloor(n-1)/2\rfloor\}$. On most of its domain, $q^{*}(n)$ is a step function with large flat regions, but it can oscillate in small transition regions. For example,
$q^{*}(n)=\begin{cases}1&\text{if }3\leq n\leq 87,\\ 2&\text{if }88\leq n\leq 224,\\ 2\text{ or }3&\text{if }225\leq n\leq 244,\\ 3&\text{if }245\leq n\leq 615,\\ 3\text{ or }4&\text{if }616\leq n\leq 645,\\ 4&\text{if }646\leq n\leq 1344,\\ \ldots\end{cases}$
where $q^{*}(225)=3$, $q^{*}(n)=2$ for $n\in\{226,\ldots,232\}$, $q^{*}(233)=3$, etc. The non-monotone ranges make it impractical to tabularize $q^{*}(n)$. Instead, we approximate $q^{*}(n)$ using a smooth function. To motivate this approximation, note that in a setting where $n$ and $q$ are large but $q$ is much smaller than $n$, which is consistent with the asymptotic regime of Assumption 2, $MSE_{j,n}(q)$ is close to $0.32q/n+0.01/q^{2}$, which is minimized at
@@
We note that the goal of this paper is to test the hypothesis $H_{0}$ that all individual hypotheses $H_{0,j}$, $j=1,\ldots,p$, hold simultaneously, while we derived the optimal $q^{*}(n)$ for an individual test statistic. Therefore, this choice does not necessarily minimize the distance between the distributions of the max-test statistic $\hat{T}\coloneqq\max_{1\leq j\leq p}\sqrt{n}\hat{\xi}_{j}$ and the bootstrap statistic $\hat{T}^{B}\coloneqq\max_{1\leq j\leq p}\hat{T}^{B}_{j}$ in any sense. However, minimizing the distance between these two distributions is considerably more difficult because the individual statistics may be arbitrarily dependent and the optimal $q$ would then depend on their (unknown) dependence structure. Developing a feasible version of this approach would require estimation of the copula of a high-dimensional random vector, which our proposal above avoids.
In simulations in Section 4, we find that our proposed choice of $q$ not only optimizes the bootstrap approximation of the individual statistics, but also yields a good bootstrap approximation for the max-statistic.

### v3

@@ Abstract
This paper proposes a nonparametric test of pairwise independence of one random variable from a large pool of other random variables. The test statistic is the maximum of several Chatterjee’s rank correlations and critical values are computed via a block multiplier bootstrap. We show in simulations that other popular tests based on distance covariances do not necessarily control size under this null. Our test, on the other hand, is shown to asymptotically control size uniformly over a large class of data-generating processes, even when the number of variables is much larger than sample size. The test is consistent against any fixed alternative. It can be combined with a stepwise procedure for selecting those variables from the pool that violate independence, while controlling the family-wise error rate. All formal results leave the dependence among variables in the pool completely unrestricted. In simulations, we find that our test is typically more powerful than competing methods (in settings where they are valid), particularly in high-dimensional scenarios or when there is dependence among variables in the pool.
Keywords: block multiplier bootstrap, Chatterjee’s rank correlation, family-wise error rate, high-dimensional data, independence test.
@@ 1 Introduction
While it has been shown that the nonparametric bootstrap does not consistently estimate the distribution of a single Chatterjee’s rank correlation (Lin and Han, 2024), the proposed block multiplier bootstrap achieves this by accounting for the dependence of the summands in the asymptotically linear representation. The existence of an asymptotic linear representation for Chatterjee’s rank correlation together with suitably strong control of the remainder terms allows us to employ recent results for the approximation of maxima of sums of high-dimensional vectors (Chernozhukov et al., 2013, 2015, 2019) to establish the validity of the block multiplier bootstrap also for the maximum of many Chatterjee’s rank correlation coefficients.
The block multiplier bootstrap requires a tuning parameter choice, namely the block size that should tend to infinity with sample size. We are able to provide a simple choice that enjoys a certain optimality property: it minimises the distance of the bootstrap variance and the asymptotic variance for each individual test statistic. In simulations, however, we find that the size and power of our test are almost insensitive to the particular value of the block size.
@@
Our paper is most closely related to the literature on testing independence of two random vectors, $H_{0}^{joint}\colon Y\perp X$, where $Y\in\mathbb{R}^{p}$ and $X\in\mathbb{R}^{q}$. This is a different hypothesis from the one we consider; for $q=1$, it implies, but is not implied by, $H_{0}$; when $H_{0}$ holds, but the copula of $Y_{1},\ldots,Y_{p}$ given $X$ is not independent of $X$, then $H_{0}^{joint}$ is violated. Therefore, tests that control the rejection probability under $H_{0}^{joint}$ are not guaranteed to control it under our hypothesis $H_{0}$. Sinha and Wieand (1977), Taskinen et al. (2005), Bakirov et al. (2006), Székely et al. (2007), Heller et al. (2012a), Heller et al. (2012b), Shi et al. (2022) propose nonparametric tests of $H_{0}^{joint}$, where $p$ and $q$ are of arbitrary, but fixed (with sample size) dimensions. Székely and Rizzo (2013) show that the test statistic of Székely et al. (2007) is biased in high dimensions and an independence test based on it therefore does not control size in high dimensions. They also propose a bias-corrected test statistic and derive its asymptotic distribution under the null of independence when the dimensions of both vectors grow with the sample size. The asymptotic regime under which their test is valid requires the dimensions of both vectors to grow, so it is not clear (at least to us) whether it is also valid when one of the two dimensions remains constant as the sample size grows. In addition, their derivation of the test statistic’s limiting distribution requires the elements of the two vectors to be exchangeable, a condition we do not require for $Y_{1},\ldots,Y_{p}$. Ramdas et al. (2015) show that both independence tests, Székely et al. (2007) and Székely and Rizzo (2013) have low power against “fair alternatives” in high dimensions. Our simulations in Section 4.4 show that these tests do not necessarily control size under $H_{0}$.
More recently, Zhou et al. (2024) and Wang et al. (2024b) propose other rank-based tests, e.g. based on Hoeffding’s D, Blum-Kiefer-Rosenblatt’s R and Bergsma-Dassios-Yanagimoto’s $\tau$ among others, of $H_{0}^{joint}$ in high dimensions. The validity of these tests relies on at least one of the dimensions $p$ and/or $q$ diverging so that a central limit theorem across the elements of, say, $Y$ can be invoked. This approach necessarily restricts the dependence of $Y_{1},\ldots,Y_{p}$, while our validity results leave the dependence completely unrestricted.
In simulations, we find that our test is typically more powerful than the alternatives by Székely et al. (2007), Székely and Rizzo (2013), Zhu et al. (2020), and Zhou et al. (2024) (in scenarios in which they are valid) in high dimensions (large $p$) or when there is dependence among the variables $Y_{1},\ldots,Y_{p}$.
Our proposed test has two additional advantages over competitors (in scenarios in which they are valid): (i) unlike tests by Zhu et al. (2020) and Zhou et al. (2024) ours is computationally inexpensive and (ii) we develop a stepwise procedure for selecting hypotheses that are not rejected so as to control the family-wise error rate.
@@
There is a large literature on nonparametric tests of mutual independence among the elements of a random vector. Some examples are Leung and Drton (2018), Yao et al. (2018), Drton et al. (2020), Wang et al. (2024a), Bastian et al. (2024); see also references therein. Xia et al. (2024) propose such a test based on Chatterjee’s rank correlation. While the hypothesis considered in our paper also involves many nonparametric independence tests, it substantially differs from the null of mutual independence. This is because, in our testing problem, $X$ occurs in every independence hypothesis and the dependence of $Y_{1},\ldots,Y_{p}$ is left unrestricted.
@@
Finally, since our proposed test is based on the rank correlation for two random variables proposed by Chatterjee (2021), our paper is also related to a recent and fast-growing literature that examines the rank correlation coefficient’s properties. Chatterjee (2021) shows asymptotic normality of the correlation coefficient under independence of the two random variables. Lin and Han (2025) and Kroll (2025) show that it is also asymptotically normal under dependence. Shi et al. (2021) and Lin and Han (2022) examine and propose improvements of the power of tests of independence based on the rank correlation coefficient. Based on the earlier work by Chatterjee (2021), Azadkia and Chatterjee (2021) introduce a graph-based correlation coefficient that can be seen as a multivariate extension of Chatterjee’s correlation coefficient. Han and Huang (2024) derive its asymptotic distribution under the null that a vector $Y$ (with fixed dimension) is independent of a random variable $X$, and they find, in simulations, that the test may be more powerful than competitors in higher dimensions. For a recent review of this literature, see Chatterjee (2024).
@@ 2 The Test
$\hat{\xi}_{j}\coloneqq 1-\frac{3n}{n^{2}-1}\sum_{i=1}^{n-1}\left\lvert\hat{F}_{Y_{j}}(Y_{j,(i+1)})-\hat{F}_{Y_{j}}(Y_{j,(i)})\right\rvert\penalty 10000\$. (1)
@@
$\xi_{j}\coloneqq\frac{\int Var(\operatorname*{\mathbb{E}}[\mathds{1}\{Y_{j}\geq t\}|X])f_{Y_{j}}(t)dt}{\int Var(\mathds{1}\{Y_{j}\geq t\})f_{Y_{j}}(t)dt},$
a measure of dependence introduced by Dette et al. (2013) in the case in which $Y_{j}$ has a continuous distribution with density $f_{Y_{j}}$. This measure has several attractive features (Chatterjee, 2021). Two features that are particularly important for the test proposed in this paper are that (i) $\xi_{j}$ is equal to zero if, and only if, $X$ and $Y_{j}$ are independent, and (ii) the estimator $\hat{\xi}_{j}$ admits an asymptotic linear representation (shown in (3) below) with a remainder that we can show to be sufficiently small. Our arguments for validity of the proposed test in high dimensions crucially depend on property (ii).
@@
$\sqrt{n}\,\hat{\xi}_{j}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}+r_{j,n}\penalty 10000\ ,$ (3)
@@
with $U_{j,i}\coloneqq F_{Y_{j}}(Y_{j,(i)})$. A naive application of the multiplier bootstrap idea would be to repeatedly draw bootstrap multipliers $\varepsilon_{1},\ldots,\varepsilon_{n}$ as independent standard normal random variables that are independent of the data $\mathbb{D}$ and then compute a critical value from the distribution of $\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}\varepsilon_{i}W_{j,i}$ given the data. However, there are two problems with this approach. First, $\{W_{j,i}\}_{i=1}^{n}$ is not an i.i.d. sequence but a 1-dependent process. Second, $\{W_{j,i}\}_{i=1}^{n}$ is not directly observed because $F_{Y_{j}}$ is unknown.
To address the first challenge, we decompose the sum $\sum_{i=1}^{n-1}W_{j,i}$ into the sums over “big” and “small” blocks formed of $\{W_{j,i}\}_{i=1}^{n}$ with the property that the big blocks are independent of each other. Formally, let $q\geq 1$ denote the size of big blocks. It is a tuning parameter to be chosen by the researcher; in Section 2.3 we develop an optimal choice of $q$. Since $\{W_{j,i}\}_{i=1}^{n}$ is a 1-dependent sequence, we let the small blocks be of length one. Further, let $m\coloneqq\lfloor(n-1)/(q+1)\rfloor$, where $\lfloor\nu\rfloor$ is the integer part of $\nu$, denote the number of big blocks. Lastly, $r\coloneqq n-1-m(q+1)$, $0\leq r<q+1$, is the number of remaining summands that are not allocated to any of the small or big blocks. With this notation, we can thus write
@@ 2.1 Size Control
Suppose $p\geq 2$. There exist constants $C_{1}>0$ and $0<\gamma<1/4$ such that $(1/q)\log^{2}p\leq C_{1}n^{-\gamma}$ and $\max\big\{q\log^{5/2}p,\,\sqrt{q}\log^{7/2}(pn)\big\}\leq C_{1}n^{1/2-\gamma}$.
@@
$T_{0}\coloneqq\max_{1\leq j\leq p}\,\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}\penalty 10000\ ,$
@@
Step (iii) of the derivation combines several results on high-dimensional Gaussian approximations from Chernozhukov et al. (2013, 2015, 2019). The main challenge of the proof is contained in steps (i) and (ii), where we need to establish that all remainder terms $r_{1,n},\ldots,r_{p,n}$ from the representation (3) vanish at a suitably fast rate. Angus (1995) shows that, for each $j$, $r_{j,n}=o_{P}(n^{-1/2})$, but in our high-dimensional setting, we need stronger control of these remainder terms, and these results are developed in Lemma 2.
@@
Remark 3 (studentisation).
The test proposed in this section does not studentise the test statistics. In simulations in Section 4, we find that studentisation may improve size and power of the test. We consider studentising the individual test statistics by their standard deviation under the null, i.e.,
@@
For the bootstrap statistics there are at least two different possibilities for studentisation. First, one could studentise it by the square root of $\operatorname*{\mathbb{E}}[A_{j,k}^{2}]/q=0.4+0.1/q$, i.e.
$\hat{T}^{B,stud\,1}\coloneqq\max_{1\leq j\leq p}\frac{1}{\sqrt{m}}\sum_{k=1}^{m}\varepsilon_{k}\frac{\hat{A}_{j,k}}{\sqrt{0.4q+0.1}}.$
This standardisation ensures that the diagonal elements of the bootstrap covariance matrix of individual tests are all approximately centred at one for any $q$, and hence approximately match the variance of $\sqrt{n}\hat{\xi}_{j}/\sqrt{v_{n}}$. The second possibility is to employ a bootstrap test statistic in which the big blocks sums are demeaned and standardised by their sample standard deviation, i.e.
$\hat{T}^{B,stud\,2}\coloneqq\max_{1\leq j\leq p}\frac{1}{\sqrt{m}}\sum_{k=1}^{m}\varepsilon_{k}\frac{\hat{A}_{j,k}-\frac{1}{m}\sum_{k=1}^{m}\hat{A}_{j,k}}{\sqrt{\frac{1}{m}\sum_{k=1}^{m}\hat{A}_{j,k}^{2}}}.$
This standardisation ensures that the diagonal elements of the bootstrap covariance matrix of individual test statistics are all equal to one for any $q$.
@@ 2.2 Consistency
In light of recent discussions related to the local asymptotic power of Chatterjee’s test of independence (e.g., Shi et al. (2021) and Lin and Han (2022)), it would be interesting to study the power of our proposed test to local alternatives. We view this analysis as beyond the scope of the paper and relegate it to future research.
@@ 2.3 Choice of $q$
In our bootstrap procedure, $T^{B}_{0,j}$ mimics the behaviour of the individual test statistic $\sqrt{n}\hat{\xi}_{j}$, which is asymptotically normally distributed and, under the null, has variance $v_{n}$ given in Remark 3. Since $q$ affects the conditional distribution of $T^{B}_{0,j}$ only through the bootstrap variance $V^{B}_{j}$ and it does not affect the test statistic itself, we aim to choose $q$ such that $V^{B}_{j}$ is close to $v_{n}$.
@@
Minimising $MSE_{j,n}(q)$ over $q\in\mathbb{N}^{+}$ yields our proposed optimal choice of $q$:
@@
The minimiser $q^{*}(n)$ can be computed for each $n$ by evaluating the function $MSE_{j,n}(q)$ over a grid of values $q\in\{1,2,\ldots,\lfloor(n-1)/2\rfloor\}$. On most of its domain, $q^{*}(n)$ is a step function with large flat regions, but it can oscillate in small transition regions. For example,
[…]
where $q^{*}(225)=3$, $q^{*}(n)=2$ for $n\in\{226,\ldots,232\}$, $q^{*}(233)=3$, etc. The non-monotone ranges make it impractical to tabularise $q^{*}(n)$. Instead, we approximate $q^{*}(n)$ using a smooth function. To motivate this approximation, note that in a setting where $n$ and $q$ are large but $q$ is much smaller than $n$, which is consistent with the asymptotic regime of Assumption 2, $MSE_{j,n}(q)$ is close to $0.32q/n+0.01/q^{2}$, which is minimised at
@@
We note that the goal of this paper is to test the hypothesis $H_{0}$ that all individual hypotheses $H_{0,j}$, $j=1,\ldots,p$, hold simultaneously, while we derived the optimal $q^{*}(n)$ for an individual test statistic. Therefore, this choice does not necessarily minimise the distance between the distributions of the max-test statistic $\hat{T}\coloneqq\max_{1\leq j\leq p}\sqrt{n}\hat{\xi}_{j}$ and the bootstrap statistic $\hat{T}^{B}\coloneqq\max_{1\leq j\leq p}\hat{T}^{B}_{j}$ in any sense. However, minimising the distance between these two distributions is considerably more difficult because the individual statistics may be arbitrarily dependent and the optimal $q$ would then depend on their (unknown) dependence structure. Developing a feasible version of this approach would require estimation of the copula of a high-dimensional random vector, which our proposal above avoids.
In simulations in Section 4, we find that our proposed choice of $q$ not only optimises the bootstrap approximation of the individual statistics, but also yields a good bootstrap approximation for the max-statistic.